Find the slope, x intercept and y intercept of the line 2/3x+y/3=-1/3
Answer by ewatrrr(24785)


Hi


2/3x+y/3=-1/3
2x + y = -1
y = -2x - 1 Slope m = -2 and the y-intercept is Pt(0,-1)
x-intercept when y = 0
0 = -2x - 1
2x = -1
x = -1/2 y-intercept is Pt(-1/2,0)